OpenStudy (anonymous):

make 'a' the subject of c=(√a^2-b^2)

OpenStudy (cwrw238):

is that the square root of a^2 only or square root of a^2 + b^2

OpenStudy (anonymous):

it's the square root of a^2 - b^2

OpenStudy (ghazi):

you mean this \[\sqrt{a^2-b^2}\] ?

OpenStudy (anonymous):

yes :)

OpenStudy (ghazi):

then just use a^2-b^2= (a-b)(a+b)

OpenStudy (anonymous):

but how do i use it in the sqrt?

hartnn (hartnn):

so eliminate sqrt by squaring both sides...then use it....

OpenStudy (ghazi):

square both the sides \[c={(a-b)(a+b)}\]

OpenStudy (ghazi):

or \[c= a^2-b^2...a^2=c-b^2\]

OpenStudy (ghazi):

now you can do it

hartnn (hartnn):

or u can do this: